OpenStudy (anonymous):

find exact value without a calculator: cot(3pi/4)

OpenStudy (debbieg):

Welcome to Open Study! :) Do you understand how to use a reference angle and the quadrant location to find trig function values for multiples of the "common angles", pi/3 (60*), pi/4 (45*), and pi/6 (30*)?

OpenStudy (debbieg):



OpenStudy (debbieg):

The function values for an angle differ from the function value for its reference angle, if at all, only in their SIGN. Since this is Q2 you know that cotangent is negative. So cot(3pi/4)=-cot(reference angle)

OpenStudy (anonymous):

Hi, I believe the exact value will be -1. I have my reasoning if you are interested. Remember OpenStudy is here for you. Have a great night! ~Chance~

OpenStudy (anonymous):

\[\cot \frac{ 3\pi }{ 4 }=\cot \left( \pi-\frac{ \pi }{4 } \right)=-\cot \frac{ \pi }{ 4 }=-1\]

OpenStudy (anonymous):

\[\cot \frac{ 3\pi }{4 }=\cot \left( \frac{ \pi }{2 }+\frac{ \pi }{4}\right)=-\tan \frac{ \pi }{4 }=-1\]

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."